OpenStudy (the_puppet1):

The variable in the equation 26 = 1 - 3r is r. Remembering the form ax + b = c, what are the values of the constants, a, b, and c in the equation? a = ___ b = ___ c = ___

OpenStudy (welshfella):

lets rearrange your equation to get it in the form ax + b = c we want the variable r on left side so firstly we add 3r to both sides 26 = 1 - 3r 26 + 3r = 1 - 3r + 3r

OpenStudy (welshfella):

3r + 26 = 1

OpenStudy (welshfella):

comparing:- ax + b = c 3r + 26 = 1

OpenStudy (welshfella):

so a = 3 what are b and c equal to?

OpenStudy (the_puppet1):

a = 3 b = 1 c = 26 Is that correct?

OpenStudy (welshfella):

no b = 26 and c = 1

OpenStudy (the_puppet1):

I won't go in on B

OpenStudy (welshfella):

???

OpenStudy (the_puppet1):

It's a = -3 b = 1 c = 26

OpenStudy (the_puppet1):

But nice try tough

OpenStudy (welshfella):

well really the 2 answers are correct because the equation can also be written as -3r + 1 = 26